The value of cos 180/11+ cos 3*180/11+cos 5*180/11+cos 7*180/11+cos 9*180/11 is



  1. 0

  2. -1/2

  3. 1/2

  4. 1


 

Grade:11

2 Answers

suryakanth AskiitiansExpert-IITB
13 years ago

Dear abhipsa,

we know that

cos a + cos 3a + cos 5a + .. + cos(2n-1)a = sin 2na/2 sin a

In our case n = 5

therefore RHS = sin 10a / (2sina)

=>RHS =  sin 2*5 *(180/11) / 2sin (180/11)

=> RHS = sin (180/11) /2 sin(180/11)

=> rhs = 1/2

hence the answer is option 3 ......1/2

GOODLUCK

Please feel free to post as many doubts on our discussion forum as you can. We are all IITians and here to help you in your IIT JEE preparation.

Win exciting gifts by answering the questions on Discussion Forum..

anil dhar
9 years ago
0